OpenStudy (anonymous):

Express the following in divided form: x ---- (x+1)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=x%2F%28x%2B1%29

jhonyy9 (jhonyy9):

x : x+1 = 1 x+1 ---- 0 +1 so x=(x+1)*1 -1